OpenStudy (anonymous):

Estimate each square root. Show work! Square Root Of 39, Help ?

myininaya (myininaya):

\[\sqrt{36}<\sqrt{39}<\sqrt{49}\]

myininaya (myininaya):

\[6<\sqrt{39}<7\]

myininaya (myininaya):

between 6 and 7 would be a good estimate

myininaya (myininaya):

39 is closer to 36 then 49 so it would be closer to 6 then it would be 7

myininaya (myininaya):

so lets just say 6.3

myininaya (myininaya):

lets see what it actually is

OpenStudy (anonymous):

we shall see myininaya!!! 6.24499... so pretty close

myininaya (myininaya):

\[\sqrt{39}\approx6.244997998\] from the calculator

OpenStudy (anonymous):

how u get the squiggly equal sign?

OpenStudy (anonymous):

Can't find it on my equation thing :(

myininaya (myininaya):

approx \[\approx \]

OpenStudy (anonymous):

so just type in approx?

myininaya (myininaya):

into the editor

OpenStudy (anonymous):

KK thanks!

myininaya (myininaya):

you see the division symbol in the editor?

OpenStudy (anonymous):

\[x \approx 3.99\]

myininaya (myininaya):

:) very good

OpenStudy (anonymous):

:)

myininaya (myininaya):

there are other ways to appoximate sqrt{39} could use linear approximation or newton's method but i think this method is pretty good( i don't know if this method has a special name or now)